# Bayes optimal learning of attention-indexed models

### v2

@@ 1 Introduction
The transformer architecture [vaswani2017attention] has transformed machine learning, achieving state-of-the-art results in natural language processing [devlin2019bertpretrainingdeepbidirectional, brown2020language], computer vision [dosovitskiy2021imageworth16x16words], and beyond. Its core innovation—the self-attention mechanism—enables models to capture long-range dependencies between tokens. Despite their empirical success, transformers remain poorly understood theoretically, especially regarding how data structure, attention bias, and training dynamics interact in finite-sample regimes. While mechanistic interpretability has shed light on trained models, the learning process itself—what is statistically and computationally learnable from limited data—remains unexplained. A common strategy toward progress is to study simplified models in high-dimensional regimes, where the blessing of dimensionality [donoho2000high] can yield tractable characterizations of learning. A key ingredient in this approach is a synthetic data model that captures salient aspects of real-world structure.
@@
Theoretical understanding of fully connected neural networks has advanced significantly through the analysis of Gaussian single-index and multi-index models in the high-dimensional limit [Arous2021, abbe22a, Veiga2022, ba2022high, arnaboldi23a, collinswoodfin2023hitting, Damian2023, bietti2023learning, moniri2023theory, berthier2024learning]. In statistical physics, similar models appear as teacher-student perceptrons [gardner1989three, sompolinsky1990learning, watkin1993, goldt2019dynamics] or committee machines [engel2001statistical, Aubin2018]. These setups typically assume i.i.d. Gaussian inputs, with targets depending on a small number of random projections—“indices”—of the input. They provide a rich theoretical playground for jointly analyzing learning dynamics, generalization, and architectural biases.
Recent work has extended this framework to model key aspects of transformers, introducing the sequence multi-index (SMI) model [cui2024phase, cui2024highdimensionallearningnarrowneural, cui_bayes-optimal_2025]. While insightful, existing SMI models require the width of the key and query matrices to be much smaller than the token embedding dimension—a regime where only narrow attention layers can be analyzed. In contrast, practical transformers typically use key and query widths comparable to the embedding dimension. This motivates our contribution: a high-dimensional yet analyzable model where learnable matrices have extensive rank. We call this the attention-indexed model.
@@
The attention-indexed model (AIM) is motivated by a generative perspective, capturing how structured token-level outputs arise from layered bilinear interactions between high-dimensional embeddings—mirroring attention computations in transformers. The idea of modeling learning through such structured synthetic data dates back to the teacher–student setting in the perceptron literature [gardner1989three, sompolinsky1990learning, engel2001statistical], and more recently to single-index and multi-index models [Arous2021, abbe22a, Veiga2022, ba2022high, arnaboldi23a, collinswoodfin2023hitting, Damian2023, bietti2023learning, moniri2023theory, berthier2024learning].
@@
While many theoretical studies explore simplified transformer variants, most do not rely on or benefit from the high-dimensional limit. These include works that analyze one-layer attention under finite embedding dimension [song2024unraveling, tian2023scan, makkuva2024local, nichani2024transformers, yangtraining], or study training dynamics in the linear, kernel, or random feature regimes [hron2020infinite, fu2023can, lu2024asymptotic]. Others use infinite-width approximations without access to generalization error [bordelon2024infinite, noci2023shaped]. By contrast, theoretical analysis of nonlinear attention layers with trainable key and query matrices in the limit of high embedding dimension—together with sharp control of generalization—is less explored. As far as we are aware, only a few works address this regime [cui2024highdimensionallearningnarrowneural, cui2024phase, troiani2025fundamentallimitslearningsequence, marion2024attention], and they all assume attention matrices of finite width.
@@
Methodologically, our approach builds on techniques from high-dimensional multi-index models, particularly those developed in [Aubin2018, troiani2024fundamental], and their recent generalizations to sequence learning with multiple low-width self-attention layers [troiani2025fundamentallimitslearningsequence]. The main technical challenge addressed in this paper is extending these tools to the case where the width $r$ of the attention matrices scales proportionally with the embedding dimension—i.e., the extensive-width regime—going beyond the key limitations of prior analyses.
@@
To tackle this, we leverage recent results on the ellipsoid fitting problem [maillard2023exact, maillard2024fitting] and its connection to two-layer neural networks with quadratic activations and extensive width [maillard_bayes-optimal_2024, xu_fundamental_2025]. Remarkably, the linear AIM model with $T=L=1$ is mathematically equivalent to such quadratic networks, allowing us to adopt these methods. We generalize this connection to arbitrary $T,L$. This is enabled by a central conceptual tool, the AIM index, which disentangles the complexity of deep attention models. It allows us to split the problem into two subproblems: (i) how structure propagates across layers and tokens, and (ii) how attention matrices are learned from those structures. This separation is crucial in extending the theory to multiple layers and tokens.
@@
Finally, we note that we focus here on the tied case $Q=K$ for clarity. The untied setting $Q\neq K$ is amenable to similar analysis following [erba2024bilinear], and we leave its treatment for future work.
@@ 2 Setting
We consider a dataset $\mathcal{D}=\{y^{\mu}\,,X_{0}^{\mu}\}$ of $n$ samples indexed by $\mu$, where $X_{0}^{\mu}=\{{\bm{x}}_{a}^{\mu\,\top}\}_{a=1}^{T}\in\mathbb{R}^{T\times d}$ has rows ${\bm{x}}_{a}^{\mu\,\top}$. Each sample consists of the embeddings of $T$ tokens ${\bm{x}}_{a}^{\mu\top}\in\mathbb{R}^{d}$, taken as standard Gaussian ${\bm{x}}_{a}^{\mu\top}\sim\mathcal{N}(0,\mathbb{I}_{d})$ and of $T\times T$ matching output matrices $y^{\mu}$ encoding pair-wise information on the original tokens. We stress that the Gaussian assumption for the data can be relaxed in the same spirit as in [xu_fundamental_2025, Assumption 2.2].
We generate $y^{\mu}$ using an attention-indexed model as given in (1) with matrices $\{S^{*}_{\ell}\}_{\ell=1,...,L}$ that are symmetric and extracted independently from a rotationally invariant ensemble $P_{S}(S)=P_{S}(O^{\top}SO)$ for any $d\times d$ rotation matrix $O$. We fix the normalizations such that $\mathbb{E}_{P_{S}}[\operatorname{Tr}S]=\kappa_{1}d$ and $\mathbb{E}_{P_{S}}[\operatorname{Tr}S^{2}]=\kappa_{2}d$ and with $\kappa_{1},\,\kappa_{2}=\mathcal{O}(1)$. We assume that the empirical spectral distribution of $S\sim P_{S}$ converges to a a distribution $\mu_{S}$ for $d\to+\infty$. This setting can be relaxed in several directions, allowing for different prior distributions $P^{(\ell)}_{S}$ for different layers, as well as considering non-symmetric matrices [erbatroiani2022].
@@ 3 Statistical and computational limits for AIMs
These expressions provide a prediction for the information-theoretically optimal estimation and generalization errors, and as such they constitute a sharp information theoretical bound on the performance of any algorithm. We derive Result 3.1 in Appendix C using the heuristic replica method, but believe that a rigorous treatment is possible by adapting [xu_fundamental_2025] to the case of multiple attention indices $L>1$ and multiple tokens $T>1$.
In our next result we provide an efficient polynomial-time approximate message passing (AMP) algorithm that in the high-dimensional limit saturates this bound under the condition that there is a unique local extremizer of (17) (if multiple extermizers are present, computational-to-statistical gaps may arise [zdeborova2016statistical]).
@@
Result 3.2 (which we derive in Appendix C) is a classic statement in the theory of AMP, that here we adapt to take into account multiple attention indices and multiple tokens. We remark that even though the denoiser $g_{\rm in}$ in (16) is a complicated non-separable function, state evolution still holds [berthier2020state, Gerbelot]. Additionally, we show in Appendix C that the fixed point of (21) are the extremisers of (17). The AMP algorithm 1 is thus both a tool that gives us theoretical guarantees though (21) and a practical algorithm that can be efficiently implemented and run on a machine. We discuss the implementation in Appendix C.
@@
where from a heavily correlated set of $L$ noisy observations $Y_{\ell}\in\mathbb{R}^{d\times d}$, $\ell=1,\dots,L$, one needs to estimate back the $L$ independent matrices $S_{m}\sim P_{S}$. Here $\Delta\in\mathbb{S}^{+}_{L}$ is an $L$-dimensional symmetric positive-definite matrix acting as a noise-to-signal ratio, and $\Xi_{\ell}$ are independent ${\rm GOE}(d)$ matrices acting as noise. To the best of our knowledge, for $L>1$ and non-diagonal $\Delta$ this is still a challenging open problem (in case of diagonal $\Delta$, the problem factorizes in $L$ independent matrix denoising problems). In practice, one can approximate rotationally-invariant matrix denoisers by low-degree spectral polynomials, following [semerjian2024matrix]; this provides a practical route to AMP for multi-layer $^{\hat{q}}$ when the exact HCIZ-based denoiser is unavailable [bun2016rotational].
For $L=1$, the problem (22) reduces to Bayes-optimal denoising of a rotationally invariant matrix [maillard_bayes-optimal_2024, erbatroiani2022]. We start by remarking that for $L=1$ the quantities $Q,q,{\hat{q}}$ governing the BO performance are scalar.
@@
where $Q=\kappa_{2}$, $\mu_{\Delta}$ is the asymptotic spectral density of a matrix $X=S+\sqrt{\Delta}Z$ with $S\sim P_{S}$ and $Z\sim{\rm GOE}(d)$, and $X=O^{\top}\Lambda O$ is the eigen-decomposition of $X$ (see [maillard_bayes-optimal_2024] for details).
@@ 4 Results for single-layer attention
Corollary 4.2 follows by combining Result 4.2 with the small-width analysis of [maillard_bayes-optimal_2024]. We remark that (30) holds for all activation (using the appropriate $g_{\rm out}$ function) and all values of $T\geq 2$. We plot the analytical prediction for the BO estimation error given in Corollary 4.2 in Figure 1 right.
@@
We plot the BO estimation error given in Result 4.3 in Figure 2 left, and observe that contrary to the hardmax case, the BO estimation error vanishes at a finite value of $\alpha$ (the so-called strong recovery threshold). Interestingly, the BO errors given in Result (4.3) is independent of the value of the inverse temperature $\beta$ and reduces to the case of a single-token model with linear activation [maillard_bayes-optimal_2024], modulo a rescaling of the sample ratio $\alpha$ to $2\alpha/(T^{2}+T-2)$ (notice that the rescaling is not just given by the total number unordered couples of tokens $T(T+1)/2$, as it would be in the case of a multi-token case with bijective activation, see App. C). The softmax activation is almost invertible, meaning that given the output, the input is fully determined apart for a common additive shift (acting as a noise correlated with the data), and is additionally constrained by the symmetry of the attention matrix. Result (4.3) precisely quantifies the amount of samples required to estimate this undetermined shift. More precisely, fix a given estimation error. Then, achieving this error with the BO estimator in the softmax case with $T\geq 2$ requires a factor $1+2/(T(T+1))$ more samples than the case of a fully bijective activation.
On the other hand, we remark that the AMP algorithm for the softmax activation at $T\geq 2$ is not a simple rescaling of the AMP for the single-token linear-activation case given in [maillard_bayes-optimal_2024]. The AMP output function $g_{\rm out}$ is given by
@@
Thanks to the mentioned reduction, one can transfer directly several results from [maillard_bayes-optimal_2024] to the case of softmax tied-attention, including an explicit prediction for the strong recovery threshold (the value of $\alpha$ after which the BO error is zero), the slope of the error at strong recovery, and the small-width and large-width limits (see App. D). In particular, the strong recovery threshold satisfies
@@
with training set generated by Eq. (27) (we optimize using the ADAM optimizer [AdamPaper]). In line with previous work [maillard_bayes-optimal_2024, erba2024bilinear], we also consider the Averaged GD (AGD) estimator given by
@@
where we average over $M$ initial matrices $W^{(0)}_{m}$, and $W^{\rm final}_{m}$ is the corresponding set of weights at convergence. We plot the results of our numerical experiments at $d=200$ for both GD and AGD in Figure 2 right. As already observed in [maillard_bayes-optimal_2024, erba2024bilinear], AGD reaches performances compatible with the BO estimation error, while GD has worse error. We remark that both variants seem to achieve perfect recovery at the BO threshold (32). This phenomenon, at this point well documented within this class of models, is still not understood.
@@ C.1 Replica analysis of AIM and their state evolution
$\quad\times\prod_{\mu=1}^{n}\Bigl[\int\!\!\prod_{u,\ell}\mathrm{d}h^{(\ell),\mu,u}_{ab}\;P_{h}\bigl(h^{(\ell),\mu,u}\bigr)\,\prod_{u,\ell,a\leq b}P_{\mathrm{\rm out}}(y^{\mu}_{ab}\mid\frac{\{h^{(\ell),\mu,u}_{ab}}{\sqrt{2-\delta_{ab}}}\}_{ab})\Bigr],$ (76)
@@ C.2 Prior Term Computation
with $Z_{\ell}$ GOE(d) matrices and $S_{\ell}\in\mathbb{R}^{d\times d}$ rotationally invariant matrices, leading to an exponential term of the form of $\frac{1}{2}\sum_{\ell}\operatorname{Tr}((\sum_{k}\sqrt{\hat{q}}_{\ell k}S_{k}-Y_{\ell})^{2})$. Such equivalence between the matrix denoising problem in (96) and (95) is analogous to those of [erba2024bilinear, maillard_bayes-optimal_2024].

### v3

@@ 1 Introduction
The transformer architecture [46] has transformed machine learning, achieving state-of-the-art results in natural language processing [26, 10], computer vision [18], and beyond. Its core innovation—the self-attention mechanism—enables models to capture long-range dependencies between tokens. Despite their empirical success, transformers remain poorly understood theoretically, especially regarding how data structure, attention bias, and training dynamics interact in finite-sample regimes. While mechanistic interpretability has shed light on trained models, the learning process itself—what is statistically and computationally learnable from limited data—remains unexplained. A common strategy toward progress is to study simplified models in high-dimensional regimes, where the blessing of dimensionality [17] can yield tractable characterizations of learning. A key ingredient in this approach is a synthetic data model that captures salient aspects of real-world structure.
@@
Theoretical understanding of fully connected neural networks has advanced significantly through the analysis of Gaussian single-index and multi-index models in the high-dimensional limit [3, 1, 47, 5, 2, 12, 16, 8, 35, 7]. In statistical physics, similar models appear as teacher-student perceptrons [22, 40, 48, 24] or committee machines [19, 4]. These setups typically assume i.i.d. Gaussian inputs, with targets depending on a small number of random projections—“indices”—of the input. They provide a rich theoretical playground for jointly analyzing learning dynamics, generalization, and architectural biases.
Recent work has extended this framework to model key aspects of transformers, introducing the sequence multi-index (SMI) model [13, 15, 14]. While insightful, existing SMI models require the width of the key and query matrices to be much smaller than the token embedding dimension—a regime where only narrow attention layers can be analyzed. In contrast, practical transformers typically use key and query widths comparable to the embedding dimension. This motivates our contribution: a high-dimensional yet analyzable model where learnable matrices have extensive rank. We call this the attention-indexed model.
@@
The attention-indexed model (AIM) is motivated by a generative perspective, capturing how structured token-level outputs arise from layered bilinear interactions between high-dimensional embeddings—mirroring attention computations in transformers. The idea of modeling learning through such structured synthetic data dates back to the teacher–student setting in the perceptron literature [22, 40, 19], and more recently to single-index and multi-index models [3, 1, 47, 5, 2, 12, 16, 8, 35, 7].
@@
While many theoretical studies explore simplified transformer variants, most do not rely on or benefit from the high-dimensional limit. These include works that analyze one-layer attention under finite embedding dimension [41, 42, 33, 36, 50], or study training dynamics in the linear, kernel, or random feature regimes [25, 21, 28]. Others use infinite-width approximations without access to generalization error [9, 38]. By contrast, theoretical analysis of nonlinear attention layers with trainable key and query matrices in the limit of high embedding dimension—together with sharp control of generalization—is less explored. As far as we are aware, only a few works address this regime [15, 13, 43, 34], and they all assume attention matrices of finite width.
@@
Methodologically, our approach builds on techniques from high-dimensional multi-index models, particularly those developed in [4, 44], and their recent generalizations to sequence learning with multiple low-width self-attention layers [43]. The main technical challenge addressed in this paper is extending these tools to the case where the width $r$ of the attention matrices scales proportionally with the embedding dimension—i.e., the extensive-width regime—going beyond the key limitations of prior analyses.
@@
To tackle this, we leverage recent results on the ellipsoid fitting problem [29, 31] and its connection to two-layer neural networks with quadratic activations and extensive width [32, 49]. Remarkably, the linear AIM model with $T=L=1$ is mathematically equivalent to such quadratic networks, allowing us to adopt these methods. We generalize this connection to arbitrary $T,L$. This is enabled by a central conceptual tool, the AIM index, which disentangles the complexity of deep attention models. It allows us to split the problem into two subproblems: (i) how structure propagates across layers and tokens, and (ii) how attention matrices are learned from those structures. This separation is crucial in extending the theory to multiple layers and tokens.
@@
Finally, we note that we focus here on the tied case $Q=K$ for clarity. The untied setting $Q\neq K$ is amenable to similar analysis following [20], and we leave its treatment for future work.
@@ 2 Setting
We consider a dataset $\mathcal{D}=\{y^{\mu}\,,X_{0}^{\mu}\}$ of $n$ samples indexed by $\mu$, where $X_{0}^{\mu}=\{{\bm{x}}_{a}^{\mu\,\top}\}_{a=1}^{T}\in\mathbb{R}^{T\times d}$ has rows ${\bm{x}}_{a}^{\mu\,\top}$. Each sample consists of the embeddings of $T$ tokens ${\bm{x}}_{a}^{\mu\top}\in\mathbb{R}^{d}$, taken as standard Gaussian ${\bm{x}}_{a}^{\mu\top}\sim\mathcal{N}(0,\mathbb{I}_{d})$ and of $T\times T$ matching output matrices $y^{\mu}$ encoding pair-wise information on the original tokens. We stress that the Gaussian assumption for the data can be relaxed in the same spirit as in [49, Assumption 2.2].
We generate $y^{\mu}$ using an attention-indexed model as given in (1) with matrices $\{S^{*}_{\ell}\}_{\ell=1,...,L}$ that are symmetric and extracted independently from a rotationally invariant ensemble $P_{S}(S)=P_{S}(O^{\top}SO)$ for any $d\times d$ rotation matrix $O$. We fix the normalizations such that $\mathbb{E}_{P_{S}}[\operatorname{Tr}S]=\kappa_{1}d$ and $\mathbb{E}_{P_{S}}[\operatorname{Tr}S^{2}]=\kappa_{2}d$ and with $\kappa_{1},\,\kappa_{2}=\mathcal{O}(1)$. We assume that the empirical spectral distribution of $S\sim P_{S}$ converges to a a distribution $\mu_{S}$ for $d\to+\infty$. This setting can be relaxed in several directions, allowing for different prior distributions $P^{(\ell)}_{S}$ for different layers, as well as considering non-symmetric matrices [45].
@@ 3 Statistical and computational limits for AIMs
These expressions provide a prediction for the information-theoretically optimal estimation and generalization errors, and as such they constitute a sharp information theoretical bound on the performance of any algorithm. We derive Result 3.1 in Appendix C using the heuristic replica method, but believe that a rigorous treatment is possible by adapting [49] to the case of multiple attention indices $L>1$ and multiple tokens $T>1$.
In our next result we provide an efficient polynomial-time approximate message passing (AMP) algorithm that in the high-dimensional limit saturates this bound under the condition that there is a unique local extremizer of (17) (if multiple extermizers are present, computational-to-statistical gaps may arise [51]).
@@
Result 3.2 (which we derive in Appendix C) is a classic statement in the theory of AMP, that here we adapt to take into account multiple attention indices and multiple tokens. We remark that even though the denoiser $g_{\rm in}$ in (16) is a complicated non-separable function, state evolution still holds [6, 23]. Additionally, we show in Appendix C that the fixed point of (21) are the extremisers of (17). The AMP algorithm 1 is thus both a tool that gives us theoretical guarantees though (21) and a practical algorithm that can be efficiently implemented and run on a machine. We discuss the implementation in Appendix C.
@@
where from a heavily correlated set of $L$ noisy observations $Y_{\ell}\in\mathbb{R}^{d\times d}$, $\ell=1,\dots,L$, one needs to estimate back the $L$ independent matrices $S_{m}\sim P_{S}$. Here $\Delta\in\mathbb{S}^{+}_{L}$ is an $L$-dimensional symmetric positive-definite matrix acting as a noise-to-signal ratio, and $\Xi_{\ell}$ are independent ${\rm GOE}(d)$ matrices acting as noise. To the best of our knowledge, for $L>1$ and non-diagonal $\Delta$ this is still a challenging open problem (in case of diagonal $\Delta$, the problem factorizes in $L$ independent matrix denoising problems). In practice, one can approximate rotationally-invariant matrix denoisers by low-degree spectral polynomials, following [39]; this provides a practical route to AMP for multi-layer $^{\hat{q}}$ when the exact HCIZ-based denoiser is unavailable [11].
For $L=1$, the problem (22) reduces to Bayes-optimal denoising of a rotationally invariant matrix [32, 45]. We start by remarking that for $L=1$ the quantities $Q,q,{\hat{q}}$ governing the BO performance are scalar.
@@
where $Q=\kappa_{2}$, $\mu_{\Delta}$ is the asymptotic spectral density of a matrix $X=S+\sqrt{\Delta}Z$ with $S\sim P_{S}$ and $Z\sim{\rm GOE}(d)$, and $X=O^{\top}\Lambda O$ is the eigen-decomposition of $X$ (see [32] for details).
@@ 4 Results for single-layer attention
Corollary 4.2 follows by combining Result 4.2 with the small-width analysis of [32]. We remark that (30) holds for all activation (using the appropriate $g_{\rm out}$ function) and all values of $T\geq 2$. We plot the analytical prediction for the BO estimation error given in Corollary 4.2 in Figure 1 right.
@@
We plot the BO estimation error given in Result 4.3 in Figure 2 left, and observe that contrary to the hardmax case, the BO estimation error vanishes at a finite value of $\alpha$ (the so-called strong recovery threshold). Interestingly, the BO errors given in Result (4.3) is independent of the value of the inverse temperature $\beta$ and reduces to the case of a single-token model with linear activation [32], modulo a rescaling of the sample ratio $\alpha$ to $2\alpha/(T^{2}+T-2)$ (notice that the rescaling is not just given by the total number unordered couples of tokens $T(T+1)/2$, as it would be in the case of a multi-token case with bijective activation, see App. C). The softmax activation is almost invertible, meaning that given the output, the input is fully determined apart for a common additive shift (acting as a noise correlated with the data), and is additionally constrained by the symmetry of the attention matrix. Result (4.3) precisely quantifies the amount of samples required to estimate this undetermined shift. More precisely, fix a given estimation error. Then, achieving this error with the BO estimator in the softmax case with $T\geq 2$ requires a factor $1+2/(T(T+1))$ more samples than the case of a fully bijective activation.
On the other hand, we remark that the AMP algorithm for the softmax activation at $T\geq 2$ is not a simple rescaling of the AMP for the single-token linear-activation case given in [32]. The AMP output function $g_{\rm out}$ is given by
@@
Thanks to the mentioned reduction, one can transfer directly several results from [32] to the case of softmax tied-attention, including an explicit prediction for the strong recovery threshold (the value of $\alpha$ after which the BO error is zero), the slope of the error at strong recovery, and the small-width and large-width limits (see App. D). In particular, the strong recovery threshold satisfies
@@
with training set generated by Eq. (27) (we optimize using the ADAM optimizer [27]). In line with previous work [32, 20], we also consider the Averaged GD (AGD) estimator given by
@@
where we average over $M$ initial matrices $W^{(0)}_{m}$, and $W^{\rm final}_{m}$ is the corresponding set of weights at convergence. We plot the results of our numerical experiments at $d=200$ for both GD and AGD in Figure 2 right. As already observed in [32, 20], AGD reaches performances compatible with the BO estimation error, while GD has worse error. We remark that both variants seem to achieve perfect recovery at the BO threshold (32). This phenomenon, at this point well documented within this class of models, is still not understood.
@@ C.1 Replica analysis of AIM and their state evolution
$\quad\times\prod_{\mu=1}^{n}\Bigl[\int\!\!\prod_{u,\ell}\mathrm{d}h^{(\ell),\mu,u}_{ab}\;P_{h}\bigl(h^{(\ell),\mu,u}\bigr)\,\prod_{u,\ell,a\leq b}P_{\mathrm{\rm out}}(y^{\mu}_{ab}\mid\{\frac{h^{(\ell),\mu,u}_{ab}}{\sqrt{2-\delta_{ab}}}\}_{ab})\Bigr],$ (76)
@@ C.2 Prior Term Computation
with $Z_{\ell}$ GOE(d) matrices and $S_{\ell}\in\mathbb{R}^{d\times d}$ rotationally invariant matrices, leading to an exponential term of the form of $\frac{1}{2}\sum_{\ell}\operatorname{Tr}((\sum_{k}\sqrt{\hat{q}}_{\ell k}S_{k}-Y_{\ell})^{2})$. Such equivalence between the matrix denoising problem in (96) and (95) is analogous to those of [20, 32].